The sum of 13 and the opposite of a number was multiplied by 3. This result was 11 less than twice the number. What was the number?
Answer by ikleyn(52777)

.
The sum of 13 and the opposite of a number was multiplied by 3.
This result was 11 less than twice the number. What was the number?
~~~~~~~~~~~~~~~~

Let x be the number, now unknown.


The problem says that

    3(13 + (-x)) = 2x - 11.


Simplify, solve and find x

    3*(13-x) = 2x - 11

    39 - 3x  = 2x - 11

    39 + 11  = 2x + 3x

       50    =    5x

        x    =    50/5 = 10.


ANSWER.  The number is 10.

Solved.